# Autalic: A Dataset for Anti-AUTistic Ableist Language In Context

### v3

@@ Abstract
As our understanding of autism and ableism continues to increase, so does our understanding of ableist language towards autistic people. Such language poses a significant challenge in NLP research due to its subtle and context-dependent nature. Yet, detecting anti-autistic ableist language remains underexplored, with existing NLP tools often failing to capture its nuanced expressions. We present Autalic, the first dataset dedicated to the detection of anti-autistic ableist language in context, addressing a significant gap in the field. Autalic comprises 2,400 autism-related sentences collected from Reddit, accompanied by surrounding context, and annotated by trained experts with backgrounds in neurodiversity. Our comprehensive evaluation reveals that current language models, including state-of-the-art LLMs, struggle both to reliably identify anti-autistic ableism and to align with human judgments, underscoring their limitations in this domain. We publicly release Autalic along with the individual annotations. This dataset serves as a crucial step towards developing more inclusive and context-aware NLP systems that better reflect diverse perspectives.
@@ 1 Introduction
There are several critical frameworks used to define autism Lawson and Beckett (2021), including the medical model, which defines disability as a “disease” and is one of the most widely used in computer science research focusing on autism Rizvi et al. (2024); Spiel et al. (2019a); Sideraki and Drigas (2021); Anagnostopoulou et al. (2020); Parsons et al. (2020); Williams et al. (2023); Sum et al. (2022).
Since this framework defines autism as a deficit of skills, its applications in technology research largely focus on providing diagnosis and treatment to autistic people Baron-Cohen (1997); Begum et al. (2016); Rizvi et al. (2024); Spiel et al. (2019b). This belief also posits neurotypical behaviors as the “norm” and autism as a “deficit” of these norms, thereby promoting neuronormativity instead of neurodiversity, which views all neurotypes as valid forms of human diversity Bottema-Beutel et al. (2021a); Walker (2014).
To improve the alignment of AI research with neurodiversity, we present Autalic, a dataset of 2,400 autism-related sentences that we collect from various communities on Reddit, along with the original context of 2,014 immediately preceding and 2,400 following sentences. We aim to fill a critical gap in current NLP research, which has largely overlooked the nuanced and context-dependent nature of ableist speech targeting autistic individuals. Our dataset not only captures key contextual elements but also incorporates a comprehensive annotation process led by trained annotators with an understanding of the autistic community, ensuring higher reliability and relevance. Our final dataset contains all of the labels to capture the nuances in human perspectives, and allows Autalic to serve as a resource for researchers studying anti-autistic ableist speech, neurodiversity, or disagreements in general.
Through a series of experiments with classical models and 4 LLMs, we find empirical evidence highlighting the difficulty of this task, and that LLMs are not reliable agents for such annotations. Our evaluations indicate that reasoning LLMs have the most consistent scores regardless of the language used in the prompt, thereby indicating a more thorough understanding of the different ways anti-autistic speech may be identified or may manifest in text. We find that in-context learning examples provide mixed results in helping improve the task comprehension among LLMs.
@@ 2 Related Work
Anti-autistic ableist language can be diverse in scope. It may include perpetuating stereotypes, using offensive language and slurs, or centering non-autistic people over the perspectives of autistic people Bottema-Beutel et al. (2021a); Rizvi et al. (2024); Darazsdi and Bialka (2023). While abusive language detection systems can help identify such speech, they are known to demonstrate bias Manerba and Tonelli (2021); Venkit et al. (2022), with even LLMs perpetuating ableist biases Gadiraju et al. (2023a). Additionally, anti-autistic ableist speech remains understudied, which is concerning given that classifiers trained on multiple hate speech datasets have shown a failure to generalize to target groups outside of the training corpus Yoder et al. (2022).
@@
Ableist language online varies, may manifest in different ways and is ever-evolving  Heung et al. (2024); Welch et al. (2023). However, toxic language datasets focusing on hate speech and abusive language have often addressed disability in general terms but have not explicitly focused on autism ElSherief et al. (2018); Ousidhoum et al. (2019). To our knowledge, there are no previous datasets specifically focused on anti-autistic speech classification, and only 3 of the 23 datasets for bias evaluation in LLMs focus on disability Gallegos et al. (2024). LLMs may be limited in that they lack an acknowledgment of context, which leads to higher rates of false positives when classifying ableist speech Phutane et al. (2024). These limitations are also found in toxicity classifiers, which excel primarily at identifying explicit ableist speech but may otherwise perpetuate harmful social biases leading to content suppression Phutane et al. (2024). Toxic language detection models, including LLM-based models, have been found to exhibit strong negative biases toward disabilities by classifying any disability-related text as toxic Narayanan Venkit et al. (2023). Further, LLMs have been observed to perpetuate implicitly ableist stereotypes Gadiraju et al. (2023b) and bias Gama (2024); Venkit et al. (2022). This, unfortunately, can sometimes be due to a research design that overlooks intra-community and disabled people’s perspectives Mondal et al. (2022), as well as autistic people’s views, which may lead to harmful stereotypes Rizvi et al. (2024); Spiel et al. (2019b). We make a step towards addressing these issues by building a dataset that focuses on ableist speech and autism by including autistic people’s perspectives during the annotation processing as recommended by Davani et al. (2023). Autalic contains all its labels and will also be useful for researchers interested in leveraging disagreements for difficult classification tasks Leonardelli et al. (2021); Pavlick and Kwiatkowski (2019).
@@ 3.1 Data Collection
We identify a methodology for curating sentences related to autism similar to prior datasets by collecting English-language sentences from Reddit  Overbay et al. (2023); Wadhwa et al. (2023); Njoo et al. (2023); Antoniak et al. (2024); Sabri et al. (2024); Allein and Moens (2024); Buz et al. (2024). The limitations of this method are detailed in Section  6.
@@ 3.1.1 Data Collection Criteria
We select Reddit as our source based on its popularity, focus on text-based content, and fewer API restrictions than X at the time of our data collection in January 2024. We search for keywords using the default search settings, which filters posts based on relevancy by prioritizing rare words in the search query, the age of the post, and the amount of likes and comments it has.  [^1] https://support.reddithelp.com/hc/en-us/articles/19695706914196-What-filters-and-sorts-are-available The search terms include “autis*”, “ASD”, “aspergers”, and “disabilit*”; the full list is available in the Appendix, Table 4.
Figure 2: An example of a sentence from our dataset. The search keyword is shown in red, while the word in blue is an example of a word defined in our glossary.
We use the identified search terms to collect the target sentence instance containing our keywords, to be labeled by the annotators, and the sentences preceding or following target instances to provide additional context. We collect 2,400 target sentences, with 2,014 preceding and 2,400 following them. Finally, we split our dataset into three parts by randomly selecting and assigning 800 unique target sentences to create three segments that were each annotated by a group of three annotators.
@@ 3.2.1 Annotator Selection
We recruit nine upper level undergraduate researchers as volunteer annotators and randomly assign them to annotate different segments of the dataset. We ensure that their involvement in our annotation process is voluntary, informed, and mutually beneficial. In particular, some participants choose to volunteer because they care deeply about the subject matter and wish to contribute to improving AI for autistic people. We select participants for whom this collaboration would provide relevant and valuable professional experience, grant them opportunities to engage in other aspects of the research, and provide mentorship and authorship recognition in accordance with the ACL guidelines.
@@
Our annotators are US-based, culturally diverse, and include people who grew up outside the United States. They are all fluent in English. Four of our annotators are gender minorities, and at least three self-identify as neurodivergent. Although we ensure the annotators were from diverse backgrounds during our recruitment process, due to the collaborative nature of our annotation process, we do not share the individual details of their identities. We also note that any personally identifiable information was destroyed upon the conclusion of our analysis and not shared outside of our research team.
@@ 3.2.3 Data Labeling
After completing the training, we assign each of the three segments of the dataset to three randomly selected annotators. Each annotator is assigned 800 unique sentences, with a goal of completing 200 annotations each week over four weeks. Annotators select from three possible labels for each sentence: “Ableist,” “Not Ableist,” or “Needs More Context.”
@@
The number of times our annotators assign each label is detailed in Table 2. To calculate our agreement scores, we consolidated labels -1 and 0 together based on feedback from our annotators that unrelated sentences needing more context could be classified as not anti-autistic in a purely granular classification.
@@
While we use the majority label as the ground truth in our analysis, in our public dataset, we will be releasing the individual labels from each annotator due to a growing interest in embracing disagreements for such classification tasks in NLP Leonardelli et al. (2021); Kralj Novak et al. (2022); Pavlick and Kwiatkowski (2019); Plank (2022); Plank et al. (2014).
Our dataset contains 2,400 sentences labeled as containing anti-autistic ableist language or not. The labels are obtained by calculating the mode from the three annotators of each data segment. Using this methodology, 242 target sentences contain examples of anti-autistic ableist language (10% of total), and 2160 sentences do not (90% of total).
@@ 3.2.4 Providing Context
While we provide additional sentences for context, the annotators are instructed to annotate the target sentence exclusively and only refer to the other sentences for additional context, such as determining whether the sentence is part of an intra-community discussion or the use of figurative speech (i.e. sarcasm). Figure 1 provides an example of a target sentence in context.
In this example, it is difficult to determine whether or not the writer had ableist intent, as it can be interpreted in multiple ways. For example, they can be critiquing the medical model, as many autistic activists do, thereby making it non-ableist. Or they could be genuinely promoting ableist misrepresentations. The contextual sentences help the annotators better understand the writer’s intent. With these sentences, it is apparent that the writer is referring to the harmful and widely discredited association of vaccines with autism, which not only promotes anti-autistic ableism in society but also puts people’s lives at risk by spreading disinformation about the benefits and harms of life-saving vaccines Gabis et al. (2022); Taylor et al. (2014); Hotez (2021).
@@ 3.2.5 Disagreements
    a tendency to use the medical model terminology or stereotypes (n=48)
@@
    a need for additional context beyond the sentences we provided
@@
Our analysis reveals a moderately strong negative correlation between the task completion time and agreement with other annotators that is statistically significant ($R=-0.644$, $p$-value: $0.0096$). This highlights the importance of our orientation as we provided it simultaneously to the annotators. The annotators who completed their task immediately after our orientation had higher agreement.
4 Experiments
4.1 Experimental Setup
We test the performance of different types of models on our dataset, including LLMs. These models are selected based on their diverse range of complexities. We also fine-tune BERT Devlin (2018) on our dataset to test classification.
4.1.1 Models
We use logistic regression (LR) with Bag of Words as the features, and using 80%-20% train-test split of the Autalic dataset.
@@ 4.1.1 Models
We also utilize fine-tuned BERT Devlin et al. (2019) as a baseline for LLM experiment results, using an 80%-20% train-test split. The F1 scores for LR and BERT (both pretrained and fine-tuned) are presented in Table 3 alongside all LLM results.
4.1.2 Prompting LLMs
With our baseline established, we use the LLMs Gemma2  Team et al. (2024), Mistral  Jiang et al. (2023), Llama3  Dubey et al. (2024), and DeepSeek  Guo et al. (2025) to classify the sentences in our dataset, and we adjust the prompts to compare each LLM’s performance. Due to limitations in computational resources, we are unable to fine-tune the LLMs.
@@ 4.1.2 Prompting LLMs
Due to the ever-evolving nature of language and variations in preferences among autistic individuals  Taboas et al. (2023), we use three different kinds of prompts to measure the consistency LLMs have in their understanding of anti-autistic ableist speech. These prompts include person-first (i.e. ’people with autism’), identity-first (i.e. ’autistic people’), and conceptual (i.e. anti-autistic) language.
@@
We include each sentence from Autalic after the aforementioned questions in our full prompt. In addition, we provide preceding and following context for each target sentence to the LLM to mimic the level of the information supplied to human annotators. We run two sets of experiments with each LLM: one that uses zero-shot prompting, and another containing engineered prompts for in-context learning verbatim from the definitions and examples provided in our annotator orientation (Appendix Section A.2).
@@
Table 3: The F1 scores of various models using person-first (PFL), identify-first (IFL), and conceptual anti-autistic (AA) prompts with and without in-context learning examples for each LLM. The best scores for each model are in bold.
@@ 4.2 Experimental Results
4.2.1 Fine-Tuning
Our experiments reveal that utilizing BERT for this classification task can lead to high rates of censorship. As BERT (unlike the other LLMs tested) is not pre-trained with instructions, we obtain its results after fine-tuning on Autalic. While the pre-trained BERT model showed poor performance indicating that it was ineffective at identifying anti-autistic ableist speech, after fine-tuning on Autalic, the model’s performance improved dramatically across all metrics, indicating that it performs better at predicting ableist speech correctly, has fewer false positives, and has a higher sensitivity to recognizing ableist speech.
4.2.2 Human-LLM Alignment
Our assessment reveals that LLMs have low levels of alignment with human perspectives and the perspectives of other LLMs, which makes them unreliable agents for such classification tasks. We assess this alignment through a measurement using Cohen’s Kappa scores. The scores shown in Figure 4 indicate the highest level of alignment was demonstrated between Gemma2 and Mistral ($k=0.34$). No LLM demonstrated alignment with our human-annotated dataset, although DeepSeek’s alignment was notably higher than the others. Overall, the LLMs had low levels of agreement with human perspectives ($M=0.091,SD=0.110$). This indicates that LLMs with less than 10 billion parameters struggle with the task of classifying anti-autistic ableist language, even when provided with in-context examples.

### v4

@@ Abstract
As our awareness of autism and ableism continues to increase, so does our understanding of ableist language towards autistic people. Such language poses a significant challenge in NLP research due to its subtle and context-dependent nature. Yet, detecting anti-autistic ableist language remains underexplored, with existing NLP tools often failing to capture its nuanced expressions. We present Autalic, the first dataset dedicated to the detection of anti-autistic ableist language in context, addressing a significant gap in the field. Autalic comprises 2,400 autism-related sentences collected from Reddit, accompanied by surrounding context, and annotated by trained experts with backgrounds in neurodiversity. Our comprehensive evaluation reveals that current language models, including state-of-the-art LLMs, struggle to reliably identify anti-autistic ableism and diverge from human judgments, underscoring their limitations in this domain. We publicly release our dataset [^1] https://nrizvi.github.io/AUTALIC.html along with the individual annotations, providing an essential resource for developing more inclusive and context-aware NLP systems that better reflect diverse perspectives.
@@ 1 Introduction
There are several critical frameworks used to define autism Lawson and Beckett (2021), including the medical model, which defines disability as a “disease” and remains one of the most widely used in computer science research focusing on autism Spiel et al. (2019a); Anagnostopoulou et al. (2020); Parsons et al. (2020); Sideraki and Drigas (2021); Sum et al. (2022); Williams et al. (2023); Rizvi et al. (2024). Since this framework defines autism as a deficit of skills, its applications in technology research largely focus on providing diagnosis and treatment to autistic people Baron-Cohen (1997); Begum et al. (2016); Spiel et al. (2019b); Rizvi et al. (2024). This belief also posits neurotypical behaviors as the “norm” and autism as a “deficit” of these norms, thereby promoting neuronormativity instead of neurodiversity, which views all neurotypes as valid forms of human diversity Bottema-Beutel et al. (2021a); Walker (2014).
To move beyond the limitations of deficit-based models and better align NLP and AI research with neurodiversity, we present Autalic, a dataset of 2,400 autism-related sentences collected from Reddit, accompanied by 2,014 preceding and 2,400 following context sentences. The dataset addresses a critical gap in current NLP research, which has largely overlooked the nuanced and context-dependent nature of ableist speech targeting autistic individuals. Autalic captures key contextual elements and includes annotations by trained experts familiar with the autistic community, ensuring higher reliability and relevance. As Autalic contains all individual labels to capture the nuances in human perspectives, it can serve as a resource for researchers studying anti-autistic ableist speech, neurodiversity, or disagreements in general.
Through a series of experiments with different baselines including 4 LLMs, our evaluation indicates that while some LLMs such as DeepSeek have the most consistent scores regardless of the language used in the prompt, thereby indicating a more thorough understanding of the different ways anti-autistic speech may be identified or may manifest in text, they remain unreliable agents for annotating such a task. We find that in-context learning examples provide mixed results in helping improve the task comprehension among LLMs. Our results provide empirical evidence highlighting the difficulty of this task.
@@ 2 Related Work
Anti-autistic ableist language can be diverse in scope. It may include perpetuating stereotypes, using offensive language and slurs, or centering non-autistic people over the perspectives of autistic people Bottema-Beutel et al. (2021a); Darazsdi and Bialka (2023); Rizvi et al. (2024). While abusive language detection systems can help identify such speech, they are known to demonstrate bias Manerba and Tonelli (2021); Venkit et al. (2022), with even LLMs perpetuating ableist biases Gadiraju et al. (2023a). Additionally, anti-autistic ableist speech remains understudied, which is concerning given that classifiers trained on multiple hate speech datasets have shown a failure to generalize to target groups outside of the training corpus Yoder et al. (2022).
@@
Ableist language online varies, may manifest in different ways and is ever-evolving  Welch et al. (2023); Heung et al. (2024). However, toxic language datasets focusing on hate speech and abusive language have often addressed disability in general terms but have not explicitly focused on autism ElSherief et al. (2018); Ousidhoum et al. (2019). To our knowledge, there are no previous datasets specifically focused on anti-autistic speech classification, and only 3 of the 23 datasets for bias evaluation in LLMs focus on disability Gallegos et al. (2024). LLMs may be limited in that they lack an acknowledgment of context, which leads to higher rates of false positives when classifying ableist speech Phutane et al. (2024). These limitations are also found in toxicity classifiers, which mainly identify explicit ableist speech but may otherwise perpetuate harmful social biases leading to content suppression Phutane et al. (2024). Toxic language detection models, including LLM-based models, have been found to exhibit strong negative biases toward disabilities by classifying any disability-related text as toxic Narayanan Venkit et al. (2023). Further, LLMs have been observed to perpetuate implicitly ableist stereotypes Gadiraju et al. (2023b) and bias Gama (2024); Venkit et al. (2022). This, unfortunately, can sometimes be due to a research design that overlooks intra-community and disabled people’s perspectives Mondal et al. (2022), as well as autistic people’s views, which may lead to harmful stereotypes Spiel et al. (2019b); Rizvi et al. (2024). We make a step towards addressing these issues by building a dataset that focuses on ableist speech and autism and including autistic people’s perspectives during the annotation process as recommended by Davani et al. (2023). Autalic contains all its labels and will also be useful for researchers interested in leveraging disagreements for difficult classification tasks Pavlick and Kwiatkowski (2019); Leonardelli et al. (2021).
@@ 3.1 Data Collection
We identify a methodology for curating sentences related to autism similar to prior datasets by collecting English-language sentences from Reddit  Overbay et al. (2023); Wadhwa et al. (2023); Njoo et al. (2023); Antoniak et al. (2024); Sabri et al. (2024); Allein and Moens (2024); Buz et al. (2024).
@@ 3.1.1 Data Collection Criteria
We selected Reddit as our data source due to its popularity, emphasis on text-based content, and fewer API restrictions compared to X (formerly Twitter) at the time of data collection in January 2024. We performed keyword searches using Reddit’s default search settings, which prioritize relevance based on factors such as the rarity of terms in the query, the post’s age, and its number of upvotes and comment.  [^2] https://support.reddithelp.com/hc/en-us/articles/19695706914196-What-filters-and-sorts-are-available The search terms include “autis*”, “ASD”, “aspergers”, and “disabilit*”; the full list is available in the Appendix, Table 4.
Figure 2: An example sentence from our dataset. The search keyword is shown in red, while the word in blue is an example of a term defined in our glossary.
We use the identified search terms to collect target sentence instances containing our keywords, which are then labeled by our annotators. To provide additional context, we also collect the sentences immediately preceding and following each target instance. In total, we collect 2,400 target sentences, along with 2,014 preceding and 2,400 following sentences. We then split the dataset into three parts by randomly selecting and assigning 800 unique target sentences to each of three annotation segments, with each segment labeled by a group of three annotators.
@@ 3.2.1 Annotator Selection
We recruit nine upper level undergraduate researchers as volunteer annotators and randomly assign them to annotate different segments of the dataset. We ensure that their involvement in our annotation process is voluntary, informed, and mutually beneficial. In particular, some participants chose to volunteer because they care deeply about the subject matter and wish to contribute to improving AI for autistic people. We select participants for whom this collaboration would provide relevant and valuable professional experience, grant them opportunities to engage in other aspects of the research, and provide mentorship and authorship recognition in accordance with the ACL guidelines.
@@
Our annotators are US-based, culturally diverse, and include people who grew up outside the US. They are all fluent in English. Four of our annotators are gender minorities, and at least three self-identify as neurodivergent. Although we ensure the annotators were from diverse backgrounds during our recruitment process, due to the collaborative nature of our annotation process, we do not share the individual details of their identities. We also note that any personally identifiable information was destroyed upon the conclusion of our analysis and not shared outside of our research team.
@@ 3.2.3 Data Labeling
After completing the training, we assign each of the three segments of the dataset to three randomly selected annotators. Each annotator is assigned 800 unique sentences, with a goal of completing 200 annotations each week over four weeks. Annotators select from three possible labels for each sentence: “Ableist”, “Not Ableist”, or “Needs More Context”.
@@
The number of times our annotators assign each label is detailed in Table 2. To calculate our agreement scores, we consolidated labels -1 and 0 together based on feedback from our annotators that unrelated sentences needing more context could be classified as not anti-autistic in a purely granular classification. While we use the majority label as the ground truth in our analysis, we release the individual labels from each annotator due to a growing interest in embracing disagreements for such classification tasks in NLP Plank et al. (2014); Pavlick and Kwiatkowski (2019); Leonardelli et al. (2021); Kralj Novak et al. (2022); Plank (2022).
@@
Our dataset contains 2,400 sentences labeled as containing anti-autistic ableist language or not. The labels are obtained by calculating the mode from the three annotators of each data segment. Using this methodology, 242 target sentences contain examples of anti-autistic ableist language (10% of total), and 2,160 sentences do not (90% of total).
@@ 3.2.4 Providing Context
While we provide additional sentences for context, the annotators are instructed to annotate the target sentence exclusively and only refer to the other sentences for additional context, such as determining whether the sentence is part of an intra-community discussion or the use of figurative speech (i.e., sarcasm). Figure 1 provides an example of a target sentence in context. That is, in this example, it is difficult to determine whether or not the writer had ableist intent, as it can be interpreted in multiple ways—they can be critiquing the medical model, as many autistic activists do, thereby making it non-ableist; or they could be genuinely promoting ableist misrepresentations.
The contextual sentences help the annotators better understand the writer’s intent. With these sentences, it is apparent that the writer is referring to the harmful and widely discredited association of vaccines with autism, which not only promotes anti-autistic ableism in society but also puts people’s lives at risk by spreading disinformation about the benefits and harms of life-saving vaccines Gabis et al. (2022); Taylor et al. (2014); Hotez (2021).
@@ 3.2.5 Disagreements
    a tendency to use the medical model terminology or stereotypes (in 48 cases),
@@
    a need for additional context beyond the sentences we provided.
@@
Our analysis reveals a moderately strong negative correlation between the task completion time and agreement with other annotators that is statistically significant (R=0.644R=-0.644, pp-value: $0.0096$). This highlights the importance of our orientation as we provided it simultaneously to the annotators. The annotators who completed their task immediately after our orientation had higher agreement.
4 Experimental Setup
We evaluate the performance of various baselines on our dataset using an 80%–20% train–test split. Additionally, we test four LLMs selected for their diverse range of model complexities.
4.1 Baselines
Logistic Regression (LR)
We use logistic regression (LR) with Bag-of-Words (BoW) features for classification.
@@ 4.1 Baselines
We fine-tune BERT Devlin (2018) on our dataset to evaluate its performance on the classification task.
4.2 Prompting LLMs
We use four LLMs: Gemma2  Team et al. (2024), Mistral  Jiang et al. (2023), Llama3  Dubey et al. (2024), and DeepSeek  Guo et al. (2025) to classify the sentences in our dataset. We adjust the prompts to compare each LLM’s performance. Due to limitations in computational resources, we are unable to fine-tune the LLMs.
@@ 4.2 Prompting LLMs
Due to the ever-evolving nature of language and variations in preferences among autistic individuals  Taboas et al. (2023), we use three different kinds of prompts to measure the consistency LLMs have in their understanding of anti-autistic ableist speech. These prompts include person-first (i.e., “people with autism”), identity-first (i.e., “autistic people”), and conceptual (i.e., “anti-autistic”) language.
@@
We include each sentence from Autalic after the aforementioned questions in our full prompt. In addition, we provide preceding and following context for each target sentence to the LLM to mimic the level of the information supplied to human annotators.
Simple vs. ICL Prompting
We run two sets of experiments with each LLM: one that uses zero-shot prompting, and another containing engineered prompts for in-context learning verbatim from the definitions and examples provided in our annotator orientation (Appendix Section A.2).
@@
Table 3: The F1 scores of various models using person-first language (PFL), identity-first language (IFL), and conceptual anti-autistic (AA) prompts with and without in-context learning examples for each LLM. The best scores for each model are in bold.
@@ 4.3 Experimental Results
4.3.1 Fine-Tuning BERT
The F1 scores for LR and BERT (both pretrained and fine-tuned) are presented in Table 3, alongside results from all LLMs. Our experiments reveal that using BERT for this classification task can initially result in high rates of censorship if deployed as-is. The pretrained BERT model performed poorly, indicating its ineffectiveness in identifying anti-autistic ableist speech. In contrast, fine-tuning BERT on Autalic led to significant performance improvements across all metrics, enabling more accurate predictions of ableist speech, fewer false positives, and greater sensitivity to subtle expressions of ableism.
4.3.2 Humans vs. LLMs
Our assessment reveals that LLMs show low levels of alignment with both human perspectives and the outputs of other LLMs, making them unreliable agents for this classification task if deployed as-is. We measure this alignment using Cohen’s Kappa scores. As shown in Figure 4, the highest level of agreement was observed between Gemma2 and Mistral ($\kappa=0.34$). No LLM demonstrated meaningful alignment with the human-annotated dataset, although DeepSeek’s agreement was notably higher than that of the other models. Overall, the LLMs exhibited low agreement with human judgments ($M=0.091$, $SD=0.110$). These results suggest that LLMs with fewer than 10 billion parameters struggle to classify anti-autistic ableist language, even when provided with in-context examples.